# Towards Adapting Federated & Quantum Machine Learning for Network Intrusion Detection: A Survey

### v1

@@ IV-B Communication Methods
Centralized Federated Learning (Client-Server Architecture): In this type of communication, there is a central server to which all the clients are connected (see Fig. 5 (a)). All the clients send their model updates to the server where the server aggregates the models received by the client to produce a global model which is then again distributed back to the clients [42].
Decentralized Federated Learning (Peer-to-Peer): This communication does not have a dedicated server for model aggregation. Clients communicate directly with each other peer-to-peer to update their models (see Fig. 5 (b)). In this setup, each client performs local training on its data, and after a certain number of training iterations, clients share their model updates with neighboring peers [37, 47]. The process continues iteratively across the network, with clients collaboratively improving the global model without relying on a centralized coordinator.
Hierarchical Federated Learning (Multi-Tier Architecture): It organizes the communication process in multiple layers, where clients in the first layer (often edge devices or local nodes) train their models on local data and then sends their model updates to higher layers for aggregation [42] (see Fig. 5 (c)). In this setup, the first layer of clients (typically referred to as leaf nodes) performs local training and sends their updates to the second layer, which consists of aggregators or intermediate nodes (such as edge servers or routers). These aggregators perform partial global aggregation of the updates received from the first layer, combining them into a more refined model. This aggregated model is then sent to the next layer for further aggregation or to a central server for the final global model.
Multi-Server Overlapping: This communication technique is used in distributed and federated learning systems, where multiple servers are involved in the communication and aggregation process [42] (see Fig. 5 (d)). In this approach, several servers, often located in different regions or networks, handle different parts of the communication and model aggregation tasks. These servers overlap in terms of the data or model updates they handle, working collaboratively rather than independently.
@@ IV-C FL Deployment
FL deployment can be categorized into cross-device and cross-silo deployment [42] (see Fig. 6).
@@ IV-D FL Model Aggregation
In FL, aggregation refers to the process of combining the model updates (usually weights or gradients) from multiple local devices to update the global model. Various aggregation techniques are used in Federated Learning methods [19, 42, 48]. This paper aims to provide a comprehensive overview of the aggregation methods, encompassing all known approaches based on our research and survey. Fig. 7 shows an overview of FL model aggregation.
@@
Federated Averaging (FedAvg): It is the most common aggregation technique in federated learning. It aggregates model updates by averaging the local model weights from each client [19]. The general formula for Federated Averaging (FedAvg) is
@@
Federated Stochastic Variance Reduced Gradient (FedSVRG): It is a variant of FedAvg that uses variance reduction to make the aggregation more efficient, aiming to reduce the number of rounds needed for convergence [39]. Each client computes a gradient approximation with reduced variance, and these approximations are aggregated similarly to FedAvg.
Federated Proximal (FedProx): This technique introduces a proximal term to the objective function, regularizing the local updates to prevent over-fitting to noisy data and reducing the effect of heterogeneous data across clients [49]. Mathematically it can be represented as:
@@
where $D_{k}$ is the dataset on client $k$, $C_{k}$ represents the local centroids, and $c_{j}$ is the $j$-th centroid. The global server aggregates the local centroids $C_{k}$ from $K$ clients to update global centroids $C$ using: $C=\frac{1}{K}\sum_{k=1}^{K}C_{k}$ [50]. This iterative process refines clustering results while ensuring that raw data remains on clients, maintaining privacy and reducing communication overhead.
Weighted Federated Averaging (WFedAvg): It is a variation of the Federated Averaging (FedAvg) algorithm, where the contribution of each client to the global model is weighted based on the size of its local dataset [42]. This ensures that clients with larger datasets have a proportionally greater influence on the aggregated global model, addressing imbalances in data distribution across clients. Mathematically, the global model update in Weighted FedAvg can be expressed as:
@@
Median-Based Federated Averaging (MFedAvg): MFedAvg is an algorithm that aggregates models using the median as a central measure. The global model update in Median Federated Averaging (MFedAvg) can be expressed as [42]:
@@
Trimmed Mean: The global model is updated by aggregating the local model updates from clients while removing extreme outliers to ensure robustness [42, 51]. The algorithm proceeds by first collecting local updates from selected clients, then sorting these updates and trimming a specified percentage of the extreme values from both ends. The remaining updates are then averaged to compute the aggregated update for the global model. Mathematically, it can be represented as:
@@
Krum(K-Center): In Krum Aggregation, the global model update is computed by selecting the local update from the client with the smallest distance-sum to its neighbors [42]. At each round, each client sends its local model update $\delta w_{i}$ to the server. The server computes the distance between every pair of updates $\delta w_{i}$ and $\delta w_{j}$ from clients $i$ and $j$, typically using Euclidean distance $d_{ij}=\|\delta w_{i}-\delta w_{j}\|$. For each client $i$, the server identifies its nearest neighbors $N_{i}$ by selecting the $n$-closest updates, where $n=p\cdot(|K|-1)$, with $p$ being the trim percentage. The client $S_{\text{min}}$ with the smallest distance sum $Dis\_Sum(i)=\sum_{j\in N_{i}}d_{ij}$ is selected, and its update $\delta w_{S_{\text{min}}}$ is used to update the global model. Mathematically, the global model update is:
@@
Stochastic Controlled Averaging for Federated Learning (SCAFFOLD): This algorithm improves federated learning by mitigating the effect of client data heterogeneity through the use of a control term $c_{i}$ for each client [52]. At each round, a subset of clients is selected, and the global model $x$ and control term $c$ are sent to them. Each client updates its local model $y_{i}$ based on the gradient of the loss function, adjusted by the difference between its control term $c_{i}$ and the global control term $c$, and computes a new control term $c_{i}^{+}$. The updates $\Delta y_{i}=y_{i}-x$ and $\Delta c_{i}=c_{i}^{+}-c_{i}$ are then communicated back to the server. The server aggregates the updates by averaging them and applies global updates to both the model parameters and control terms, adjusting the global model $x$ using $\eta_{g}\Delta x$ and the global control term $c$ using $\frac{|S|}{N}\Delta c$, where $|S|$ is the number of clients involved in the round and $N$ is the total number of clients. This process helps to ensure more stable convergence in federated learning with non-i.i.d. data.
@@ IV-E FL Privacy Preservation
The following section delves into the diverse privacy preservation techniques utilized in Federated Learning, highlighting their role in safeguarding data privacy while enabling collaborative model training [37, 53, 20].
Homomorphic Encryption: It is a form of encryption [54] that allows computations to be performed on cipher texts (encrypted data) without the need to decrypt them first. This is powerful because it enables sensitive data to remain encrypted while being processed, ensuring privacy and security throughout the computation process [42]. The result of the computation, when decrypted, matches the result that would have been obtained if the operations had been performed on the plain text data. For example, in additive homomorphic encryption, $E(m1)+E(m2)=E(m1+m2)$, where $E(.)$ represents the encryption operation [55].
Differential Privacy: It is a privacy-preserving technique that ensures individual data points in a dataset cannot be identified or inferred from the outputs of a data analysis or computation. The main idea is to add noise to the data or query results in a way that provides useful aggregate insights while protecting individual privacy [42]. One common method to achieve differential privacy is the Laplace mechanism, which adds noise drawn from the Laplace distribution to the output of a query. If $\mathcal{A}$ is a query function (e.g., counting or summing data), the Laplace mechanism perturbs the result by adding Laplace noise scaled by $\frac{\Delta f}{\epsilon}$, where $\Delta f$ is the sensitivity of the function (the maximum possible change in the output caused by a single data point change) and is [56]:
@@
Secure Multiparty Computation: It is a cryptographic technique that allows multiple parties to jointly compute a function over their private inputs without revealing those inputs to each other [42]. The core idea is to divide the computation into smaller parts and distribute them among the parties, where each party only performs its part of the computation and shares the results in such a way that the final outcome can be obtained, but no party learns anything about the other parties’ private inputs [57]. Mathematically, let $f(x_{1},x_{2},\dots,x_{n})$ be the function to be computed, where $x_{i}$ represents the private input of party $i$. In SMPC, each party computes a share $s_{i}$ of the input and sends it to the other parties. The final result $f(x_{1},x_{2},\dots,x_{n})$ is then computed through the secure aggregation of the shares, typically via secret sharing or garbled circuits. For example, using additive secret sharing, the inputs $x_{i}$ are split into shares $s_{ij}$ such that $x_{i}=\sum_{j}s_{ij}$, and the parties collaboratively compute the result by combining their shares while maintaining the privacy of each $x_{i}$.
@@ IV-F FL Network Protocols
Communication is one of the most important parts of FL. There can be various issues related to privacy and traffic flow [58] while communicating in FL. This led to the development of various network protocols particularly focusing on FL. Network protocols are mainly focused on wireless communication which is necessary for various industries such as healthcare [59]. In this paper, we review several network protocols and techniques designed to enhance the efficiency, privacy, and robustness of Federated Learning systems [35].
@@ IV-G FL Frameworks
FL frameworks are software libraries or platforms developed to implement FL. These frameworks provide tools, APIs, and prebuilt functionalities to enable efficient communication, training, and aggregation across distributed clients in an FL setup [53]. The following frameworks have been widely adopted in research and industry for implementing FL-based methods:
  • TensorFlow Federated (TFF): It is an open-source FL framework developed by Google for experimenting and building FL algorithms [60]. It comes with a built-in simulation environment that allows developers to simulate federated learning scenarios before deploying them to real devices. It has pre-built components for FL research, and customizable algorithms and allows integration with TensorFlow [53].
  • PySyft: It is an open-source library developed by OpenMined for secure and private machine learning, including FL [61]. Its main focus is on privacy-preserving techniques in FL such as differential privacy and secure multi-party computation [53]. The core component of a PySyft library is the tensor which follows PyTorch or TensorFlow APIs but provides functionalities like remote execution and encryption.
  • Federated AI Tech Enabler (FATE): It is developed by WeBank and its primary focus is on industrial applications of FL [62]. It supports different types of FL such as horizontal, vertical and transfer FL. It also has built-in modules for security [53].
  • Flower: It is a lightweight and modular FL framework for research and production [63]. It has cross-framework compatibility (TensorFlow, PyTorch), supports edge devices and customizable strategies [63].
  • OpenFL: It is developed by Intel and is an open-source FL framework for collaborative machine learning across multiple organizations [64, 65]. Its primary focus is on cross-silo FL [66], and secure data collaboration. It also supports various ML frameworks.
  • PaddleFL: This FL framework is based on PaddlePaddle by Baidu [67] with an stan 2.0 license. It supports horizontal and vertical FL, encryption and is designed for large-scale industrial applications [53]. This package includes FedAvg, SecAgg, and differentially private stochastic gradient descent (DPSGD).
  • FedML: It is an open FL library for real-world FL and edge AI [68]. It supports hierarchical FL, edge device integration, and benchmarking tools [68].
  • LEAF: It is a benchmark framework designed for evaluating FL algorithms on realistic datasets and workloads [69]. Its primary focus is on non-IID data, customizable workloads, and extensibility [69].
@@
  • FedScale: It is an benchmarking FL framework that is designed to scale FL to large numbers of clients [71, 72]. FedScale primary focus is on scalability, optimized resource management, and dynamic scheduling of FL tasks [71].
  • Fed-TGAN: It is an FL framework for TGANs (Tabular Generative Adversarial Networks) that utilizes GANs for synthetic data generation in FL settings. It combines FL with GANs for data synthesis and improves FL for non-IID and limited datasets [73].
  • FLUTE (Federated Learning Utilities for Testing and Experimentation): It is an open-source, high-performance framework developed by Microsoft Research for simulating federated learning scenarios at scale [74].
  • IBM FL: It is an open-source framework developed by IBM that enables collaborative machine learning across organizations without sharing raw data. It focuses on secure cross-silo training and supports multiple machine learning algorithms, including both classical and deep learning models [75].
  • IDSoft: An FL framework focusing on intelligent distributed systems and soft computing [76]. It combines FL with distributed soft computing methods for resource optimization and adaptive learning [76].
Table III gives a comparison between various open-source FL frameworks [77].
@@ IV-H Model Compression Techniques for Communication Efficient Federated Learning
FL is used in various sectors such as healthcare, smart agriculture and finance [78]. One of the major challenges in FL is communicating model updates to the central server for aggregation. The complexity and size of the models can greatly impact the FL process. This highlights the need for efficient communication strategies [79] such as various model compression techniques to reduce the communication overhead without compromising the model’s performance.
@@
The most common types of model compression techniques include quantization, pruning, and knowledge distillation [13]. Most of the existing work on FL uses full-precision weights that need to be transmitted to the server for aggregation. Quantization reduces memory and computation requirements by representing model parameters such as weights and activations using fewer bits compared to standard full-precision floating point numbers [80, 81]. This greatly reduces the size and computation requirements of the model without affecting the accuracy. Oh et al. in [82] and Oh et al. in [83] presented novel techniques FedQCS (Federated Learning via Quantized Compressed Sensing) and FedVQCS (Federated Learning via Vector Quantized Compressed Sensing) respectively. In the FedQCS method, the gradients are compressed in the edge devices sequentially by block sparsification, dimension reduction and quantization. The compressed gradients are reconstructed in parameter server (PS) by minimum mean square error approach (MMSE) using expectation-maximization generalized-approximate-message-passing (EM-GAMP) algorithm [84]. The convergence rate of the proposed method is given by $O\left(\frac{1}{\sqrt{T}}\right)$, where T is the number of total iterations of the SGD algorithm. The results demonstrated that FedQCS with one bit overhead per gradient entry can attain a classification accuracy similar to the perfect reconstruction without compression. It outperformed other existing QCS-based frameworks both in terms of classification accuracy and normalized MSE of the gradient reconstruction. Similarly, in the FedVQCS method, the local model is compressed using dimensionality reduction followed by vector quantization. Gradients are reconstructed on the parameter server (PS) using a sparse signal recovery algorithm, in contrast to MMSE in [82]. The results showed that it achieved a 2.4% improvement in accuracy compared to other state-of-the-art FL methods in the MNIST and FEMNIST datasets when the communication overhead is 0.1 bits per model entry.
@@
Most of the participating clients in the FL process involve devices having resource constraints such as limited power and computational and communication capabilities. The participating clients with weak capabilities can become the bottleneck of model training. To address these issues, model pruning is one of the effective ways. It reduces the redundant or less important parameters of a neural network which helps in reducing the size of the model, computational complexity and storage costs [85]. Jiang et al. in [86] developed Federated Model Pruning (FedMP) that simultaneously helps in reducing communication and computation overhead by adaptive model pruning over heterogeneous participating clients. A Multi-Armed Bandit (MAB) was implemented to adaptively determine pruning ratios for the heterogeneous participating clients. The parameter server (PS) determines different pruning ratios for different heterogeneous clients. A novel parameter synchronization scheme Residual Recovery Synchronous Parallel (R2SP) was also implemented to ensure the sub-models’ parameters are properly synchronized before model aggregation. The proposed framework was implemented on a physical platform and the results showed that this method was efficient for various heterogeneous scenarios and can provide up to 4.1x speedup as compared to other existing FL solutions. Similarly, Yi et al. in [87] developed FedPE, a communication-efficient FL framework that allows each participating client to select personalized optimal local subnets for each round of FL. In each federated cycle, the server sends personalized subnets to each participating client. The clients test these subnets on their respective local data and compare the accuracies from previous round accuracies. If there are $k$ clients and $t$ represents each federated cycle, the difference between the accuracies at $t$ and $t+1$ is given by:
@@
Similarly, knowledge distillation is one of the methods of model compression in which knowledge is transferred from a larger mentor model to a smaller mentee model which reduces the communication overhead while maintaining high accuracy [88, 89]. Wu et al. in [90] proposed a novel technique called FedKD, a communication-efficient FL framework that communicates model parameters from a small mentee model in contrast to a large full-size model in standard FL solutions. It significantly reduces communication overhead while maintaining high accuracy. To maintain the performance of the mode, an adaptive mutual distillation strategy is introduced, where the mentor and mentee models learn reciprocally. Additionally, an adaptive loss weighting mechanism ensures high-quality knowledge transfer, and a Singular Value Decomposition (SVD)–based compression technique further reduces model update sizes. The results demonstrated a reduction in communication costs by 94.98% outperforming other FL solutions in non-IID settings. Chen et al. in [91] proposed a resource-aware transfer knowledge-based FL framework for resource-constrained devices. The proposed framework implements two grouping strategies, random and computing capacity grouping strategies. In random grouping, participating clients are grouped arbitrarily without the need for any private information whereas in computing capacity grouping clients are grouped on the basis for the computing capacity of the clients. Participating clients with limited computing capabilities are assigned lightweight models without overloading weaker devices. If the clients are participating in a first learning task, it either initializes the blank model, or in the case of subgroup training, the student model is adopted and the joint loss function is applied to incorporate knowledge distillation. The joint loss function is
@@
Table IV provides a comparison of common model compression techniques used in machine learning and federated learning for network security applications [92]. It summarizes the working principle, key advantages, and potential limitations of Quantization, Pruning, and Knowledge Distillation, which are widely adopted to reduce model size, improve inference speed, and maintain performance in resource-constrained environments.

### v2

@@ IV-B Communication Methods
Centralized Federated Learning (Client-Server Architecture): In this type of communication, there is a central server to which all the clients are connected (see Fig. 5 (a)). All the clients send their model updates to the server where the server aggregates the models received by the client to produce a global model which is then again distributed back to the clients [khraisat2024survey].
Decentralized Federated Learning (Peer-to-Peer): This communication does not have a dedicated server for model aggregation. Clients communicate directly with each other peer-to-peer to update their models (see Fig. 5 (b)). In this setup, each client performs local training on its data, and after a certain number of training iterations, clients share their model updates with neighboring peers [mothukuri2021survey, yuan2024decentralized]. The process continues iteratively across the network, with clients collaboratively improving the global model without relying on a centralized coordinator.
Hierarchical Federated Learning (Multi-Tier Architecture): It organizes the communication process in multiple layers, where clients in the first layer (often edge devices or local nodes) train their models on local data and then sends their model updates to higher layers for aggregation [khraisat2024survey] (see Fig. 5 (c)). In this setup, the first layer of clients (typically referred to as leaf nodes) performs local training and sends their updates to the second layer, which consists of aggregators or intermediate nodes (such as edge servers or routers). These aggregators perform partial global aggregation of the updates received from the first layer, combining them into a more refined model. This aggregated model is then sent to the next layer for further aggregation or to a central server for the final global model.
Multi-Server Overlapping: This communication technique is used in distributed and federated learning systems, where multiple servers are involved in the communication and aggregation process [khraisat2024survey] (see Fig. 5 (d)). In this approach, several servers, often located in different regions or networks, handle different parts of the communication and model aggregation tasks. These servers overlap in terms of the data or model updates they handle, working collaboratively rather than independently.
@@ IV-C FL Deployment
FL deployment can be categorized into cross-device and cross-silo deployment [khraisat2024survey] (see Fig. 6).
@@ IV-D FL Model Aggregation
In FL, aggregation refers to the process of combining the model updates (usually weights or gradients) from multiple local devices to update the global model. Various aggregation techniques are used in Federated Learning methods [qi2024model, khraisat2024survey, nanayakkara2024understanding]. This paper aims to provide a comprehensive overview of the aggregation methods, encompassing all known approaches based on our research and survey. Fig. 7 shows an overview of FL model aggregation.
@@
Federated Averaging (FedAvg): It is the most common aggregation technique in federated learning. It aggregates model updates by averaging the local model weights from each client [qi2024model]. The general formula for Federated Averaging (FedAvg) is
@@
Federated Stochastic Variance Reduced Gradient (FedSVRG): It is a variant of FedAvg that uses variance reduction to make the aggregation more efficient, aiming to reduce the number of rounds needed for convergence [li2021survey]. Each client computes a gradient approximation with reduced variance, and these approximations are aggregated similarly to FedAvg.
Federated Proximal (FedProx): This technique introduces a proximal term to the objective function, regularizing the local updates to prevent over-fitting to noisy data and reducing the effect of heterogeneous data across clients [li2020federated]. Mathematically it can be represented as:
@@
where $D_{k}$ is the dataset on client $k$, $C_{k}$ represents the local centroids, and $c_{j}$ is the $j$-th centroid. The global server aggregates the local centroids $C_{k}$ from $K$ clients to update global centroids $C$ using: $C=\frac{1}{K}\sum_{k=1}^{K}C_{k}$ [islam2024fedclust]. This iterative process refines clustering results while ensuring that raw data remains on clients, maintaining privacy and reducing communication overhead.
Weighted Federated Averaging (WFedAvg): It is a variation of the Federated Averaging (FedAvg) algorithm, where the contribution of each client to the global model is weighted based on the size of its local dataset [khraisat2024survey]. This ensures that clients with larger datasets have a proportionally greater influence on the aggregated global model, addressing imbalances in data distribution across clients. Mathematically, the global model update in Weighted FedAvg can be expressed as:
@@
Median-Based Federated Averaging (MFedAvg): MFedAvg is an algorithm that aggregates models using the median as a central measure. The global model update in Median Federated Averaging (MFedAvg) can be expressed as [khraisat2024survey]:
@@
Trimmed Mean: The global model is updated by aggregating the local model updates from clients while removing extreme outliers to ensure robustness [khraisat2024survey, 10200010]. The algorithm proceeds by first collecting local updates from selected clients, then sorting these updates and trimming a specified percentage of the extreme values from both ends. The remaining updates are then averaged to compute the aggregated update for the global model. Mathematically, it can be represented as:
@@
Krum(K-Center): In Krum Aggregation, the global model update is computed by selecting the local update from the client with the smallest distance-sum to its neighbors [khraisat2024survey]. At each round, each client sends its local model update $\delta w_{i}$ to the server. The server computes the distance between every pair of updates $\delta w_{i}$ and $\delta w_{j}$ from clients $i$ and $j$, typically using Euclidean distance $d_{ij}=\|\delta w_{i}-\delta w_{j}\|$. For each client $i$, the server identifies its nearest neighbors $N_{i}$ by selecting the $n$-closest updates, where $n=p\cdot(|K|-1)$, with $p$ being the trim percentage. The client $S_{\text{min}}$ with the smallest distance sum $Dis\_Sum(i)=\sum_{j\in N_{i}}d_{ij}$ is selected, and its update $\delta w_{S_{\text{min}}}$ is used to update the global model. Mathematically, the global model update is:
@@
Stochastic Controlled Averaging for Federated Learning (SCAFFOLD): This algorithm improves federated learning by mitigating the effect of client data heterogeneity through the use of a control term $c_{i}$ for each client [praneeth2019scaffold]. At each round, a subset of clients is selected, and the global model $x$ and control term $c$ are sent to them. Each client updates its local model $y_{i}$ based on the gradient of the loss function, adjusted by the difference between its control term $c_{i}$ and the global control term $c$, and computes a new control term $c_{i}^{+}$. The updates $\Delta y_{i}=y_{i}-x$ and $\Delta c_{i}=c_{i}^{+}-c_{i}$ are then communicated back to the server. The server aggregates the updates by averaging them and applies global updates to both the model parameters and control terms, adjusting the global model $x$ using $\eta_{g}\Delta x$ and the global control term $c$ using $\frac{|S|}{N}\Delta c$, where $|S|$ is the number of clients involved in the round and $N$ is the total number of clients. This process helps to ensure more stable convergence in federated learning with non-i.i.d. data.
@@ IV-E FL Privacy Preservation
The following section delves into the diverse privacy preservation techniques utilized in Federated Learning, highlighting their role in safeguarding data privacy while enabling collaborative model training [mothukuri2021survey, kholod2020open, zhang2021survey].
Homomorphic Encryption: It is a form of encryption [rivest1978data] that allows computations to be performed on cipher texts (encrypted data) without the need to decrypt them first. This is powerful because it enables sensitive data to remain encrypted while being processed, ensuring privacy and security throughout the computation process [khraisat2024survey]. The result of the computation, when decrypted, matches the result that would have been obtained if the operations had been performed on the plain text data. For example, in additive homomorphic encryption, $E(m1)+E(m2)=E(m1+m2)$, where $E(.)$ represents the encryption operation [li2018cryptographic].
Differential Privacy: It is a privacy-preserving technique that ensures individual data points in a dataset cannot be identified or inferred from the outputs of a data analysis or computation. The main idea is to add noise to the data or query results in a way that provides useful aggregate insights while protecting individual privacy [khraisat2024survey]. One common method to achieve differential privacy is the Laplace mechanism, which adds noise drawn from the Laplace distribution to the output of a query. If $\mathcal{A}$ is a query function (e.g., counting or summing data), the Laplace mechanism perturbs the result by adding Laplace noise scaled by $\frac{\Delta f}{\epsilon}$, where $\Delta f$ is the sensitivity of the function (the maximum possible change in the output caused by a single data point change) and is [9714350]:
@@
Secure Multiparty Computation: It is a cryptographic technique that allows multiple parties to jointly compute a function over their private inputs without revealing those inputs to each other [khraisat2024survey]. The core idea is to divide the computation into smaller parts and distribute them among the parties, where each party only performs its part of the computation and shares the results in such a way that the final outcome can be obtained, but no party learns anything about the other parties’ private inputs [li2020privacy]. Mathematically, let $f(x_{1},x_{2},\dots,x_{n})$ be the function to be computed, where $x_{i}$ represents the private input of party $i$. In SMPC, each party computes a share $s_{i}$ of the input and sends it to the other parties. The final result $f(x_{1},x_{2},\dots,x_{n})$ is then computed through the secure aggregation of the shares, typically via secret sharing or garbled circuits. For example, using additive secret sharing, the inputs $x_{i}$ are split into shares $s_{ij}$ such that $x_{i}=\sum_{j}s_{ij}$, and the parties collaboratively compute the result by combining their shares while maintaining the privacy of each $x_{i}$.
@@ IV-F FL Network Protocols
Communication is one of the most important parts of FL. There can be various issues related to privacy and traffic flow [choudhury2002anonymizing] while communicating in FL. This led to the development of various network protocols particularly focusing on FL. Network protocols are mainly focused on wireless communication which is necessary for various industries such as healthcare [8859260]. In this paper, we review several network protocols and techniques designed to enhance the efficiency, privacy, and robustness of Federated Learning systems [9153560].
@@ IV-G FL Frameworks
FL frameworks are software libraries or platforms developed to implement FL. These frameworks provide tools, APIs, and prebuilt functionalities to enable efficient communication, training, and aggregation across distributed clients in an FL setup [kholod2020open]. The following frameworks have been widely adopted in research and industry for implementing FL-based methods:
  • TensorFlow Federated (TFF): It is an open-source FL framework developed by Google for experimenting and building FL algorithms [TFF]. It comes with a built-in simulation environment that allows developers to simulate federated learning scenarios before deploying them to real devices. It has pre-built components for FL research, and customizable algorithms and allows integration with TensorFlow [kholod2020open].
  • PySyft: It is an open-source library developed by OpenMined for secure and private machine learning, including FL [ziller2021pysyft]. Its main focus is on privacy-preserving techniques in FL such as differential privacy and secure multi-party computation [kholod2020open]. The core component of a PySyft library is the tensor which follows PyTorch or TensorFlow APIs but provides functionalities like remote execution and encryption.
  • Federated AI Tech Enabler (FATE): It is developed by WeBank and its primary focus is on industrial applications of FL [FATE]. It supports different types of FL such as horizontal, vertical and transfer FL. It also has built-in modules for security [kholod2020open].
  • Flower: It is a lightweight and modular FL framework for research and production [beutel2020flower]. It has cross-framework compatibility (TensorFlow, PyTorch), supports edge devices and customizable strategies [beutel2020flower].
  • OpenFL: It is developed by Intel and is an open-source FL framework for collaborative machine learning across multiple organizations [reina2021openfl, OpenFL]. Its primary focus is on cross-silo FL [mittone2024pushing], and secure data collaboration. It also supports various ML frameworks.
  • PaddleFL: This FL framework is based on PaddlePaddle by Baidu [BaiduFL] with an stan 2.0 license. It supports horizontal and vertical FL, encryption and is designed for large-scale industrial applications [kholod2020open]. This package includes FedAvg, SecAgg, and differentially private stochastic gradient descent (DPSGD).
  • FedML: It is an open FL library for real-world FL and edge AI [FedML]. It supports hierarchical FL, edge device integration, and benchmarking tools [FedML].
  • LEAF: It is a benchmark framework designed for evaluating FL algorithms on realistic datasets and workloads [caldas2018leaf]. Its primary focus is on non-IID data, customizable workloads, and extensibility [caldas2018leaf].
@@
  • FedScale: It is an benchmarking FL framework that is designed to scale FL to large numbers of clients [lai2022fedscale, FedScale]. FedScale primary focus is on scalability, optimized resource management, and dynamic scheduling of FL tasks [lai2022fedscale].
  • Fed-TGAN: It is an FL framework for TGANs (Tabular Generative Adversarial Networks) that utilizes GANs for synthetic data generation in FL settings. It combines FL with GANs for data synthesis and improves FL for non-IID and limited datasets [zhao2021fed].
  • FLUTE (Federated Learning Utilities for Testing and Experimentation): It is an open-source, high-performance framework developed by Microsoft Research for simulating federated learning scenarios at scale [garcia2022flute].
  • IBM FL: It is an open-source framework developed by IBM that enables collaborative machine learning across organizations without sharing raw data. It focuses on secure cross-silo training and supports multiple machine learning algorithms, including both classical and deep learning models [ludwig2020ibm].
  • IDSoft: An FL framework focusing on intelligent distributed systems and soft computing [alotaibi2023idsoft]. It combines FL with distributed soft computing methods for resource optimization and adaptive learning [alotaibi2023idsoft].
Table III gives a comparison between various open-source FL frameworks [riedel2024comparative].
@@ IV-H Model Compression Techniques for Communication Efficient Federated Learning
FL is used in various sectors such as healthcare, smart agriculture and finance [kairouz2021advances]. One of the major challenges in FL is communicating model updates to the central server for aggregation. The complexity and size of the models can greatly impact the FL process. This highlights the need for efficient communication strategies [le2022insights] such as various model compression techniques to reduce the communication overhead without compromising the model’s performance.
@@
The most common types of model compression techniques include quantization, pruning, and knowledge distillation [le2024survey]. Most of the existing work on FL uses full-precision weights that need to be transmitted to the server for aggregation. Quantization reduces memory and computation requirements by representing model parameters such as weights and activations using fewer bits compared to standard full-precision floating point numbers [liu2021zero, cai2020zeroq]. This greatly reduces the size and computation requirements of the model without affecting the accuracy. Oh et al. in [oh2022communication] and Oh et al. in [oh2023fedvqcs] presented novel techniques FedQCS (Federated Learning via Quantized Compressed Sensing) and FedVQCS (Federated Learning via Vector Quantized Compressed Sensing) respectively. In the FedQCS method, the gradients are compressed in the edge devices sequentially by block sparsification, dimension reduction and quantization. The compressed gradients are reconstructed in parameter server (PS) by minimum mean square error approach (MMSE) using expectation-maximization generalized-approximate-message-passing (EM-GAMP) algorithm [vila2011expectation]. The convergence rate of the proposed method is given by $O\left(\frac{1}{\sqrt{T}}\right)$, where T is the number of total iterations of the SGD algorithm. The results demonstrated that FedQCS with one bit overhead per gradient entry can attain a classification accuracy similar to the perfect reconstruction without compression. It outperformed other existing QCS-based frameworks both in terms of classification accuracy and normalized MSE of the gradient reconstruction. Similarly, in the FedVQCS method, the local model is compressed using dimensionality reduction followed by vector quantization. Gradients are reconstructed on the parameter server (PS) using a sparse signal recovery algorithm, in contrast to MMSE in [oh2022communication]. The results showed that it achieved a 2.4% improvement in accuracy compared to other state-of-the-art FL methods in the MNIST and FEMNIST datasets when the communication overhead is 0.1 bits per model entry.
@@
Most of the participating clients in the FL process involve devices having resource constraints such as limited power and computational and communication capabilities. The participating clients with weak capabilities can become the bottleneck of model training. To address these issues, model pruning is one of the effective ways. It reduces the redundant or less important parameters of a neural network which helps in reducing the size of the model, computational complexity and storage costs [10330640]. Jiang et al. in [jiang2023computation] developed Federated Model Pruning (FedMP) that simultaneously helps in reducing communication and computation overhead by adaptive model pruning over heterogeneous participating clients. A Multi-Armed Bandit (MAB) was implemented to adaptively determine pruning ratios for the heterogeneous participating clients. The parameter server (PS) determines different pruning ratios for different heterogeneous clients. A novel parameter synchronization scheme Residual Recovery Synchronous Parallel (R2SP) was also implemented to ensure the sub-models’ parameters are properly synchronized before model aggregation. The proposed framework was implemented on a physical platform and the results showed that this method was efficient for various heterogeneous scenarios and can provide up to 4.1x speedup as compared to other existing FL solutions. Similarly, Yi et al. in [yi2024fedpe] developed FedPE, a communication-efficient FL framework that allows each participating client to select personalized optimal local subnets for each round of FL. In each federated cycle, the server sends personalized subnets to each participating client. The clients test these subnets on their respective local data and compare the accuracies from previous round accuracies. If there are $k$ clients and $t$ represents each federated cycle, the difference between the accuracies at $t$ and $t+1$ is given by:
@@
Similarly, knowledge distillation is one of the methods of model compression in which knowledge is transferred from a larger mentor model to a smaller mentee model which reduces the communication overhead while maintaining high accuracy [hong2022analysis, ge2024pfl]. Wu et al. in [wu2022communication] proposed a novel technique called FedKD, a communication-efficient FL framework that communicates model parameters from a small mentee model in contrast to a large full-size model in standard FL solutions. It significantly reduces communication overhead while maintaining high accuracy. To maintain the performance of the mode, an adaptive mutual distillation strategy is introduced, where the mentor and mentee models learn reciprocally. Additionally, an adaptive loss weighting mechanism ensures high-quality knowledge transfer, and a Singular Value Decomposition (SVD)–based compression technique further reduces model update sizes. The results demonstrated a reduction in communication costs by 94.98% outperforming other FL solutions in non-IID settings. Chen et al. in [10090471] proposed a resource-aware transfer knowledge-based FL framework for resource-constrained devices. The proposed framework implements two grouping strategies, random and computing capacity grouping strategies. In random grouping, participating clients are grouped arbitrarily without the need for any private information whereas in computing capacity grouping clients are grouped on the basis for the computing capacity of the clients. Participating clients with limited computing capabilities are assigned lightweight models without overloading weaker devices. If the clients are participating in a first learning task, it either initializes the blank model, or in the case of subgroup training, the student model is adopted and the joint loss function is applied to incorporate knowledge distillation. The joint loss function is
@@
Table IV provides a comparison of common model compression techniques used in machine learning and federated learning for network security applications [dantas2024comprehensive]. It summarizes the working principle, key advantages, and potential limitations of Quantization, Pruning, and Knowledge Distillation, which are widely adopted to reduce model size, improve inference speed, and maintain performance in resource-constrained environments.